OpenStudy (anonymous):

Find the value of y so that the line passing through (2, 6) and (1, y) has a slope of 5.

OpenStudy (anonymous):

\[\frac{y-6}{1-2}=5\] \[\frac{y-6}{-1}=5\] \[y-6=-5\]